# Debatable Intelligence: Benchmarking LLM Judges via Debate Speech Evaluation

### v1

@@ Abstract
We introduce Debate Speech Evaluation as a novel and challenging benchmark for assessing LLM judges. Evaluating debate speeches requires a deep understanding of the speech at multiple levels, including argument strength and relevance, the coherence and organization of the speech, the appropriateness of its style and tone, and so on. This task involves a unique set of cognitive abilities that have previously received limited attention in systematic LLM benchmarking. To explore such skills, we leverage a dataset of over $600$ meticulously annotated debate speeches and present the first in-depth analysis of how state-of-the-art LLMs compare to human judges on this task. Our findings reveal a nuanced picture: while larger models can approximate individual human judgments in some respects, they differ substantially in their overall judgment behavior. We also investigate the ability of frontier LLMs to generate persuasive, opinionated speeches, showing that models may perform at a human level on this task.
@@ 1 Introduction
  • Q1: How do LLM judges perform in the task of debate speech evaluation?
@@ 2 Related Work
Multiple works in the field of computational argumentation Wachsmuth et al. (2017) have explored the quality of individual arguments Toledo et al. (2019); Gretz et al. (2020); Durmus et al. (2024); Habernal and Gurevych (2016). With the increased long-context capabilities of LLMs, researchers have begun exploring lengthy argumentative texts, applying LLMs to evaluate entire debates. A notable thread of research is using LLMs to determine the winner in a debate between two sides. Rescala et al. (2024) provide debater exchanges to a model and prompts it to predict the victor. Liang et al. (2024) introduce a more sophisticated architecture that addresses long context length challenges in multi-round debates. Their model evaluates the debate speech by speech, using analysis of previous speeches rather than the entire debate history. Liu et al. (2024) analyze how several LLMs (GPT-3.5, GPT-4, and Llama-2-70B) predict debate winners and study their biases.
@@ 3 Speech Quality Dataset
Each speech in the dataset was carefully rated by 15 experienced human annotators. Given the speech and the topic of the debate, the annotators were asked to score the following statement on a Likert scale between 1 and 5: “This speech is a good opening speech to support the topic”, as presented in Figure 1. Slonim et al. further employed additional measures to ensure the annotation quality, such as “control” speeches added to identify unreliable annotators.
In this study, we use $631$ speeches from this dataset (see Appendix A). Table 1 outlines the properties of the data.
@@
The different sources exhibit varying levels of “artificiality”: speeches of some sources (e.g., Speech-GPT2) are entirely artificial, while others are partly artificial (e.g., Arg-Human1,2), alongside $152$ completely human-authored speeches. This diversity gives the data a rich variety of writing styles and qualities. We provide additional details regarding the different sources in Appendix A.
(a) Average pair agreement
(b) Agreement with average human score
@@ 4 Experimental Setup
We run a series of experiments to assess the performance of judges (Q1) and assess their behavior (Q2). These experiments involve running multiple LLMs as judges on the speech dataset and measuring their agreement with humans. Finally, we use GPT-4.1 to generate speeches, which we then evaluate with the highest-performing judges (Q3).
@@
Our experiments encompass $23$ models of different sizes, families, and licensing types (both proprietary and open), including state-of-the-art generative and reasoning models. We use various sizes of Qwen-2.5, Llama-3.1, Llama-3.2, and Llama-3.3. Additionally, we incorporate Claude-3.5, GPT-4o, and GPT-4.1 models. We also employ four reasoning models: Claude-3.7-Sonnet, o3, o3-mini, and o4-mini. Appendix B describes additional details regarding execution.
@@
The prompt for the judge models follows the instructions given to human annotators in Slonim et al. (2021). It asks the model to score input speeches on a scale of 1 to 5. We also create a variant of this prompt that requires judges to justify their scores. Prompts are provided in Appendix B.
@@
We define the following metrics for measuring agreement on speech ratings. Following Slonim et al. (2021), we employ weighted Kappa agreement. For human annotators, we identify pairs who share at least $50$ speeches (termed the “minimal-sample” threshold) and calculate their average pairwise agreement scores. We use a leave-one-out approach to evaluate LLM judges: for each human pair, we substitute one annotator with an LLM and measure agreement on the same speech set. We perform this substitution for both positions in each pair and average the pairwise agreements.
@@
Even when enforcing a minimal-sample value, the number of shared speeches per annotator pair remains small, which could introduce noise. We therefore employ an additional, more robust metric that measures ranking correlation between an LLM judge and the average of 15 human annotations per speech, over the entire set of speeches. We use the Kendall’s Tau-C correlation metric as it can handle agreement between variables with different possible values – in our case, a continuous variable (the average human scores) and a categorical one (the LLM judge scores).
@@
We use GPT-4.1 to generate new speeches. We request the model to generate a 600-word speech about a given debate topic. Appendix C shows the prompt used for this step, along with additional technical details regarding maintaining the length limit of the speeches.
@@ 5.1 Judge Performance Evaluation
We start by analyzing the judges’ performance in Figure 2(a), aiming to examine to what extent the judges are comparable with human annotators, and can replace them on this task. It presents the average pairwise Kappa agreement between humans, along with results from repeating the experiment while replacing one annotator with various LLM judges (as described in Section 4). Our results indicate that there are LLM judges who perform on par with humans, with some, like Qwen-72B, even surpassing human performance.
@@
Figure 2(a) shows that models larger than 7B are generally more capable on this task. However, smaller models such as Qwen-7B and Qwen-14B outperform considerably stronger and more advanced models like GPT-4o. To address potential noise from the limited number of speeches shared between annotator pairs, we introduce a more robust analysis. As described in Section 4, we compute the agreement between average human scores per speech and LLM judges using Kendall’s Tau-C.
@@
Enhancing zero-shot judges
Inspired by recent works Liu et al. (2023); Luo et al. (2023), we test whether Chain-of-Thought (CoT) prompting can help to improve the performance of LLM judges. To this end, we repeat the previous experiment using a CoT variant of our prompt (available in Appendix B). Figure 3 presents our results. We observe that for most larger models, CoT prompting offers a slight improvement. Interestingly, it degrades some of the smaller judges. We hypothesize that some performance decrease could be attributed to increased parsing issues (see Appendix D for further analysis).
@@
Figure 5: Judge ratings of different speech sources, compared to human ratings (rightmost panels). Text in gray shows the Pearson correlation of human and LLMs mean scores across sources. As models become more capable, they rank speeches from different sources more similarly to how humans do. Interestingly, while stronger models align with humans on the relative ranking of speeches, they tend to assign lower overall scores to synthetic data.
@@ 5.2 Judge Behavior Analysis
The score distribution helps explain why some LLMs outperform others in one metric but not in another. We observe that models whose score distributions resemble those of humans generally obtain better Kappa scores, even if their agreement on the speech ranking is relatively low. For example, Llama-3.1-8B has a higher Kappa agreement than Llama-3.1-70B, but a lower Kendall’s tau score. This could be attributed to Kappa normalizing agreement based on the expected agreement by chance, which is derived from the scores distribution. This is in contrast to Kendall’s tau, which is a non-parametric measure of agreement.
@@
Our findings could be explained by stronger judges showing a more decisive behavior, amplifying the gap between strong and weak “systems” (in our case, groups of speeches; see Gera et al., 2024). Another factor that might contribute to the difference between human and LLM judges’ behaviour is that, in contrast to LLMs, multiple humans annotate the same speech, smoothing out extremes and biases of individual judges.
@@ 5.3 Speech Quality of Modern LLMs
Having established that certain LLM judges strongly agree with humans, we examine how the best-performing judges from each family (those with the highest Tau-C correlation) evaluate speeches generated by a SOTA model. We use GPT-4.1 to generate $152$ speeches (two per debate topic, similarly to speeches by humans) and compare their average rating to that of human speeches. The generation prompt is provided in Appendix C.
Our results, presented in Figure 6, indicate a strong preference for GPT-4.1 speeches over those authored by humans. This is in stark contrast to the results in Section 5.2, where older synthetic sources were given a very low rating. All five models, including those not subject to self-bias like Claude-3.7 and Llama-3.3-70B, rated the generated speeches higher than the human-authored ones. While we lack human annotations for further verification, these results demonstrate the rapid progress of generative models in recent years. More importantly, the ability of these models to argue controversial topics more effectively than human experts raises concerns about their safe deployment and potential misuse. We provide an example of a GPT-4.1 generated speech in Appendix H, Figure 19.
@@ 5.4 Judge Reasoning Analysis
In addition, the key points indicate that GPT-4.1 speeches are the most coherent and clear. This is unsurprising, considering how the speeches were constructed. For example, Arg-Human1 speeches consist of automatically concatenated arguments supporting the topic, while human speeches are transcribed from recordings. In Appendix H, we provide examples of multiple speeches that show the far more constructed and organized GPT-4.1 speeches. The most notable difference is the division into paragraphs and a clear structure, including an opening, three arguments, and a summary. Given that LLM judges are notoriously known to be biased towards stylistic properties Wu and Aji (2023), this may be one of the reasons that speeches by GPT-4.1 receive higher scores.
@@ 6 Discussion
The comparison of LLM judges to high-quality human annotations reveals some complex patterns. On the one hand, some LLMs are on par with individual humans in terms of instance-level agreement, occasionally even surpassing the performance of the typical human annotator. On the other, our deeper analysis demonstrates that all judges differ starkly from humans in their judgment behaviors – LLM judges assign lower evaluation scores overall, and draw sharper distinctions between “strong” and “weak” speeches, potentially exaggerating performance gaps in ways that differ from human annotators. Moreover, judges still fail to accurately replicate the system-level ranking of speech sources given by humans. These patterns echo recent works on the distinctions between system-level and instance-level judgment performance Gera et al. (2024); Gao et al. (2024). Importantly, they point to open questions around the design of LLM-as-a-Judge evaluation pipelines and whether current methods adequately reflect nuanced human judgments.
@@
Improving the judging capabilities of smaller models remains an important goal. Further research is also needed to better align LLM judgment behavior with human evaluations, particularly in capturing nuanced patterns and system-level preferences. Finally, future work may explore enabling LLM judges to rate distinct aspects of speech quality, such as rhetorical style or coherence, rather than relying solely on an overall score.
@@ 7 Conclusion
In this work, we benchmark LLMs using data collected by Slonim et al. (2021). This imposes some limitations on our experiments. First, multiple quality dimensions, such as relevance, style, factuality, and argument strength, are all combined into a single score. This may introduce interpretability challenges, which we aim to address in Section 5.4. Second, the data by Slonim et al. (2021) does not contain annotations for speeches by current state-of-the-art models. We try to address this gap in Section 5.3 by generating additional speeches using the state-of-the-art GPT-4.1. Moreover, analyzing how LLM judges rate synthetic outputs from older and newer models offers valuable insight into the qualities they prioritize in a speech, as seen in Section 5.4.
Finally, in line with the data, we focus solely on opening debate speeches. This neglects the complexities of evaluating subsequent turns in debate, where arguments and counterarguments build upon each other. Evaluating LLMs’ ability to judge multi-turn debate dynamics remains an open challenge.
@@
In Section 5.3, we use five high-performing judges to analyze speeches authored by GPT-4.1. While our experiments show these models align well with human judges, we acknowledge that actual human annotation of this specific data could yield different results. Specifically, as we point out in Section 5.4, stylistic bias Wu and Aji (2023) may have contributed to the higher score of these speeches. In adddition, judges from the GPT-4.1 series may be biased towards self-generated texts Xu et al. (2024).
@@
We make our data and code publicly available to ensure long-term reproducibility and enable others to build upon our work. The released speeches were collected by Slonim et al. (2021), which examines ethical concerns regarding the data annotation. Our analysis reveals potential risks of LLM misuse, though we recognize that studying these risks poses its own ethical challenges. For example, our findings could be used to identify models that excel at generating persuasive text beyond human capabilities and misuse them. This underscores the critical need for robust system safeguards and ethical deployment guidelines to prevent dangerous use. We used AI assistants for grammatical corrections and code writing (e.g., GitHub Copilot).

### v2

@@ Abstract
We introduce Debate Speech Evaluation as a novel and challenging benchmark for assessing LLM judges. Evaluating debate speeches requires a deep understanding of the speech at multiple levels, including argument strength and relevance, the coherence and organization of the speech, the appropriateness of its style and tone, and so on. This task involves a unique set of cognitive abilities that previously received limited attention in systematic LLM benchmarking. To explore such skills, we leverage a dataset of over $600$ meticulously annotated debate speeches and present the first in-depth analysis of how state-of-the-art LLMs compare to human judges on this task. Our findings reveal a nuanced picture: while larger models can approximate individual human judgments in some respects, they differ substantially in their overall judgment behavior. We also investigate the ability of frontier LLMs to generate persuasive, opinionated speeches, showing that models may perform at a human level on this task.
@@ 1 Introduction
  • Q1: How well do LLM judges perform the task of debate speech evaluation?
@@ 2 Related Work
Multiple works in the field of computational argumentation Wachsmuth et al. (2017) have explored the quality of individual arguments Toledo et al. (2019); Gretz et al. (2020); Durmus et al. (2024); Habernal and Gurevych (2016). With the increased long-context capabilities of LLMs, researchers have begun exploring lengthy argumentative texts, applying LLMs to evaluate entire debates. A notable thread of research is using LLMs to determine the winner in a debate between two sides. Rescala et al. (2024) provide debater exchanges to a model and prompt it to predict the victor. Liang et al. (2024) introduce a more sophisticated architecture that addresses long context length challenges in multi-round debates. Their model evaluates the debate speech by speech, using analysis of previous speeches rather than the entire debate history. Liu et al. (2024) analyze how several LLMs (GPT-3.5, GPT-4, and Llama-2-70B) predict debate winners and study their biases.
@@ 3 Speech Quality Dataset
Each speech in the dataset was carefully rated by multiple crowd annotators. Given the speech and the topic of the debate, the annotators were asked to score the following statement on a Likert scale between $1$ and $5$: “This speech is a good opening speech to support the topic”, as presented in Figure 1. To ensure annotation quality, Slonim et al. further employed additional measures, such as “control” speeches added to identify unreliable annotators. This process yielded a total of $15$ trusted human annotations per speech. Appendix C.1 provides further details on the annotation protocol and quality control measures used in Slonim et al. (2021).
In this study, we analyze $631$ speeches (see Appendix C.2) drawn from the dataset introduced by Slonim et al. (2021). Table 1 outlines the main properties of the dataset.
@@
The different sources exhibit varying levels of “artificiality”: speeches of some sources (e.g., Speech-GPT2) are entirely artificial, while others are partly artificial (e.g., Arg-Human1,2), alongside $152$ completely human-authored speeches. This diversity gives the data a rich variety of writing styles and qualities. We provide additional details regarding the different sources in Appendix C.1.
@@ 4 Experimental Setup
We run a series of experiments to assess the performance of judges (Q1) and assess their behavior (Q2). These experiments involve running multiple LLMs as judges on the speech dataset
and measuring their agreement with humans.
Finally, we use GPT-4.1 to generate speeches, which we then evaluate with the highest-performing judges (Q3).
(a) Average pair agreement
(b) Ranking agreement with average human scores
@@
Our experiments encompass $23$ models of different sizes, families, and licensing types (both proprietary and open), including state-of-the-art generative and reasoning models. We use various sizes of Qwen-2.5, Llama-3.1, Llama-3.2, and Llama-3.3. Additionally, we incorporate Claude-3.5, GPT-4o, and GPT-4.1 models. We also employ four reasoning models: Claude-3.7-Sonnet, o3, o3-mini, and o4-mini. Appendix A describes additional details regarding execution.
@@
The prompt for the judge models follows the instructions given to human annotators in Slonim et al. (2021). It asks the model to score input speeches on a scale of 1 to 5. We also create a variant of this prompt that requires judges to justify their scores. Prompts are provided in Appendix A.
@@
We define the following metrics for measuring agreement on speech ratings. Following Slonim et al. (2021), we employ weighted Kappa agreement. For human annotators, we identify pairs who share at least $50$ speeches (termed the “minimal-sample” threshold) and calculate their average pairwise agreement scores. This threshold ensures that agreement is estimated on a sufficiently robust sample, reducing the variance that would arise from very small overlaps.
We adopt a leave-one-out evaluation setup, which is common in the LLM-as-a-Judge literature Zheng et al. (2023). For each pair of human annotators who rated the same set of speeches, we replace one annotator with an LLM and compute agreement with the remaining human on the overlapping subset. We perform this substitution for both positions in each pair and average the pairwise agreements.
@@
Even when enforcing a minimal-sample value, the number of shared speeches per annotator pair remains limited, which could introduce noise and leaves much of the dataset unused. We therefore employ an additional, more robust metric that measures ranking correlation between an LLM judge and the average of $15$ human annotations per speech, over the entire set of speeches. We use the Kendall’s Tau-C correlation metric as it can handle agreement between variables with different possible values – in our case, a continuous variable (the average human scores) and a categorical one (the LLM judge scores).
@@
We use GPT-4.1 to generate new speeches. We request the model to generate two $600$-word speeches about a given debate topic (similarly to human-authored speeches). Appendix B shows the prompt used for this step, along with additional technical details.
@@ 5.1 Judge Performance Evaluation
We start by analyzing the judges’ performance, aiming to examine to what extent the judges are comparable with human annotators, and can replace them on this task. Figure 2(a) presents the average pairwise Kappa agreement between humans, along with results from repeating the experiment while replacing one annotator with various LLM judges (as described in Section 4). Our results indicate that there are LLM judges who perform on par with humans, with some, like Qwen-72B, even surpassing human performance — meaning that their agreement with a human annotator is higher than human agreement.
@@
Figure 2(a) shows that models larger than 7B are generally more capable on this task. However, smaller models such as Qwen-7B and Qwen-14B seem to outperform considerably stronger and more advanced models like GPT-4o. To address potential noise from the limited number of speeches shared between annotator pairs, we introduce a more robust analysis. As described in Section 4, we compute the ranking agreement between average human scores per speech and LLM judges using Kendall’s Tau-C. This score represents the ability of a judge to reproduce the relative ranking of speeches given by the human annotators.
@@
Figure 5: Judge ratings of different speech sources, compared to human ratings (rightmost panels). Text in gray shows the Pearson correlation between the mean scores per source given by humans and LLMs. As models become more capable, they rank speeches from different sources more similarly to how humans do. Interestingly, while stronger models align with humans on the relative ranking of speeches, they tend to assign lower overall scores to synthetic data.
Enhancing zero-shot judges
Inspired by recent works Liu et al. (2023); Luo et al. (2023), we test whether Chain-of-Thought (CoT) prompting can help to improve the performance of LLM judges. To this end, we repeat the previous experiment using a CoT variant of our prompt (available in Appendix A). Figure 3 presents our results. For most larger models, CoT prompting offers a slight improvement. Interestingly, it degrades some of the smaller judges. We hypothesize that some performance decrease could be attributed to increased parsing issues (see Appendix D).
@@ 5.2 Judge Behavior Analysis
The score distribution also helps explain why some LLMs outperform others in one metric but not in another. We observe that models whose score distributions resemble those of humans generally obtain better Kappa scores, even if their agreement on the speech ranking is relatively low. For example, Llama-3.1-8B has a higher Kappa agreement than Llama-3.1-70B, but a lower Kendall’s tau score. This could be attributed to Kappa normalizing agreement based on the expected agreement by chance, which is derived from the scores distribution. This is in contrast to Kendall’s tau, which is a non-parametric measure of agreement.
@@
Our findings could be explained by stronger judges showing a more decisive behavior, amplifying the gap between strong and weak “systems” (in our case, groups of speeches; see Gera et al., 2024). Another factor that might contribute to the difference between human and LLM judges’ behaviour is that, in contrast to LLMs, multiple humans annotate the same speech, smoothing out extremes and possible biases of individual judges.
@@ 5.3 Speech Quality of Modern LLMs
Having established that certain LLM judges strongly agree with humans, we examine how the best-performing judges from each family (those with the highest Tau-C correlation) evaluate speeches generated by a SOTA model. We use GPT-4.1 to generate $152$ speeches (two per debate topic, similarly to speeches by humans) and compare their average rating to that of human speeches. The generation prompt is provided in Appendix B.
Our results, presented in Figure 6, indicate a strong preference for GPT-4.1 speeches over those authored by humans. This is in stark contrast to the results in Section 5.2, where older synthetic sources were given a very low rating.
All five models, including those not subject to self-bias like Claude-3.7 and Llama-3.3-70B, rated the generated speeches higher than the human-authored ones. While we lack human annotations for further verification, these results demonstrate the rapid progress of generative models in recent years. More importantly, the ability of these models to argue controversial topics – ostensibly, in a more effective manner than human experts – raises concerns about their safe deployment and potential misuse. We provide an example of a GPT-4.1 generated speech in Appendix H, Figure 19.
@@ 5.4 Judge Reasoning Analysis
In addition, the key points indicate that GPT-4.1 speeches are the most coherent and clear. This is unsurprising, considering how the speeches were constructed. For example, Arg-Human1 speeches consist of automatically concatenated arguments supporting the topic, while human speeches are transcribed from recordings. The speech examples we provide in Appendix H illustrate the far more constructed and organized nature of
GPT-4.1 speeches. The most notable difference is the division into paragraphs and a clear structure, including an opening, three arguments, and a summary. Given that LLM judges are notoriously
biased towards stylistic properties Wu and Aji (2023), this may be one of the reasons that speeches by GPT-4.1 receive higher scores.
@@ 6 Discussion
The comparison of LLM judges to high-quality human annotations reveals some complex patterns. On the one hand, some LLMs are on par with individual humans in terms of instance-level agreement, occasionally even surpassing the performance of the typical human annotator. On the other, our deeper analysis demonstrates that all judges differ starkly from humans in their judgement behaviors – LLM judges assign lower evaluation scores overall, and draw sharper distinctions between “strong” and “weak” speeches, potentially exaggerating performance gaps in ways that differ from human annotators. Moreover, judges still fail to accurately replicate the system-level ranking of speech sources given by humans. These patterns echo recent works on the distinctions between system-level and instance-level judgment performance Gera et al. (2024); Gao et al. (2024). Importantly, they point to open questions around the design of LLM-as-a-Judge evaluation pipelines and whether current methods adequately reflect nuanced human judgments.
@@
Improving the judging capabilities of smaller models remains an important goal, as this would expand accessibility and reduce the computational cost of deploying LLM-as-a-Judge systems. Beyond model size, further research is needed to better align LLM judgment behavior with human evaluations, particularly in capturing nuanced patterns and system-level preferences.
Another promising direction is to explore multi-dimensional evaluations, where judges provide separate ratings for aspects such as rhetorical style, coherence, factual grounding, and persuasiveness, complementing the holistic score. Our preliminary use of Key Point Analysis (Section 5.4) illustrates the potential of this approach and shows how richer frameworks can surface more fine-grained distinctions in model behavior.
Future work may also examine the role of prompt design more systematically. Here we focused on a controlled zero-shot setting to isolate inherent judgment abilities. However, in-context learning (ICL) setups, where few-shot demonstrations are included, may substantially impact performance, particularly for weaker models. A broader comparison between zero-shot, CoT, and ICL conditions could help disentangle genuine judging capabilities from adaptation to prompt structure.
Finally, future work may investigate temporal alignment between training data, human annotations, and generated content. The speeches in our dataset are from 2020 or earlier, whereas modern LLMs are trained on more recent corpora. Although our topics are timeless, many domains evolve quickly, raising the question of whether improved speech quality reflects better reasoning, updated knowledge, or both. Systematic studies of knowledge and discourse shifts over time are needed to answer such questions.
@@ 7 Conclusion
In this work, we benchmark LLMs using data collected by Slonim et al. (2021). This imposes some limitations on our experiments. First, multiple quality dimensions, such as relevance, style, factuality, and argument strength, are all combined into a single score. This may introduce interpretability challenges, which we aim to address in Section 5.4. Second, the data by Slonim et al. (2021) does not contain annotations for speeches by current SOTA models.
We try to address this gap in Section 5.3 by generating additional speeches using the state-of-the-art GPT-4.1. Moreover, analyzing how LLM judges rate synthetic outputs from older and newer models offers valuable insight into the qualities they prioritize in a speech, as seen in Section 5.4.
Finally, in line with the data, we focus solely on opening debate speeches. This neglects the complexities of evaluating subsequent turns in debate, where arguments and counterarguments build upon each other. Evaluating LLMs’ ability to judge multi-turn dynamics remains an open challenge.
@@
In Section 5.3, we use five high-performing judges to analyze speeches authored by GPT-4.1. While our experiments show these models align well with human judges, we acknowledge that actual human annotation of this specific data could yield different results. Specifically, as we point out in Section 5.4, stylistic bias Wu and Aji (2023) may have contributed to the higher score of these speeches. In addition, judges from the GPT-4.1 series may be biased toward self-generated texts Xu et al. (2024).
Prompt sensitivity
While we used the same prompt structure across models to maintain a controlled comparison setting, we recognize that this can disadvantage some models and benefit others. Our choice reflects a trade-off between per-model optimization and fair benchmarking. Importantly, optimizing prompts separately for each model risks entangling model-specific tuning with the evaluation itself, making it harder to draw generalizable conclusions. We acknowledge, however, that examining the sensitivity of results to alternative prompts is valuable, and we leave a systematic investigation of this aspect for future work.
Inference reproducibility
For efficiency, we ran models larger than 88B parameters using 4-bit quantization, as full-precision inference was not feasible under our resource constraints. While quantization may slightly affect performance, we applied it consistently across all relevant models to ensure fairness.
@@
We make our data and code publicly available to ensure long-term reproducibility and enable others to build upon our work. The released speeches were collected by Slonim et al. (2021), which examines ethical concerns regarding the data annotation. We note that some debate topics in this dataset involve sensitive issues, and certain speeches could be considered offensive to some readers.
Our analysis also reveals potential risks of LLM misuse, though we recognize that studying these risks poses its own ethical challenges. For example, our findings could be used to identify models that excel at generating persuasive text beyond human capabilities and misuse them. This underscores the critical need for robust system safeguards and ethical deployment guidelines to prevent dangerous use.